An angle measures 12° less than the measure of its supplementary angle. What is the measure of each angle?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39626)
Answer by ikleyn(52858)

.
An angle measures 12° less than the measure of its supplementary angle. What is the measure of each angle?
~~~~~~~~~~~~~~~~

Let x be the angle measure.

Then the measure of other angle is (x+12) degrees, according to the condition.


Since the angles are supplementary, they sum up to 180 degrees

    
    x + (x+12) = 180


Simplify and solve


    2x = 180 - 12 = 168

     x            = 168/2 = 84.


ANSWER.  The angle measure is 84°.  The other angle is  84° + 12° = 96°.

Solved.

----------------

To learn about supplementary angles,  read the lessons
    - Angles basics
    - HOW TO solve problems on supplementary, complementary or vertical angles - Examples
in this site.


Consider these lessons as your handbook,  textbook,  guide,  tutorials and  (free of charge)  home teacher.